# Testing Low-Resource Language Support in LLMs Using Language Proficiency Exams: the Case of Luxembourgish

### v2

@@ F Fine-grained Exam Results
Qwen 2 0.5 9.5 26.2 22.7 34.8 16.7 16.2 20.0 21.8 0.0 24.0 23.8 30.8 21.4 14.3 25.9 20.5 41.4 0.0 1.5
Llama 3.2 1 19.0 15.5 6.8 0.0 0.0 0.0 0.0 3.6 0.0 4.0 9.5 0.0 0.0 0.0 9.3 0.0 2.0 0.0 0.0
Qwen 2 1.5 33.3 39.3 27.3 30.4 16.7 37.8 55.0 20.0 22.2 20.0 28.6 23.1 0.0 0.0 40.7 35.9 38.2 9.1 13.6
StableLM 2 1.6 23.8 16.7 18.2 4.3 16.7 8.1 5.0 12.7 11.1 14.0 9.5 15.4 14.3 14.3 18.5 10.3 4.6 0.0 4.5
Gemma 2 2 28.6 31.0 20.5 43.5 33.3 29.7 25.0 23.6 22.2 24.0 28.6 7.7 35.7 0.0 24.1 20.5 40.1 18.2 19.7
Llama 3.2 3 28.6 35.7 29.5 34.8 33.3 18.9 15.0 27.3 44.4 24.0 19.0 38.5 21.4 14.3 35.2 20.5 39.5 22.7 30.3
Phi 3 3.8 23.8 23.8 36.4 21.7 33.3 32.4 35.0 23.6 11.1 26.0 19.0 30.8 21.4 42.9 24.1 25.6 38.2 29.5 33.3
Phi 3.5 3.8 38.1 19.0 11.4 8.7 33.3 2.7 5.0 5.5 11.1 4.0 4.8 0.0 7.1 0.0 3.7 5.1 9.2 13.6 10.6
Qwen 2 7 33.3 39.3 31.8 21.7 50.0 40.5 45.0 25.5 33.3 28.0 23.8 30.8 21.4 42.9 27.8 48.7 48.0 40.9 37.9
Llama 2 7 19.0 16.7 6.8 8.7 33.3 0.0 0.0 7.3 11.1 8.0 4.8 0.0 21.4 0.0 9.3 0.0 31.6 9.1 7.6
WizardLM 2 7 23.8 21.4 25.0 8.7 33.3 32.4 15.0 25.5 33.3 24.0 33.3 15.4 35.7 28.6 25.9 28.2 33.6 20.5 25.8
Mistral 7 28.6 36.9 25.0 17.4 16.7 29.7 35.0 29.1 22.2 26.0 38.1 15.4 14.3 28.6 20.4 33.3 40.8 25.0 37.9
Aya-23 8 28.6 29.8 27.3 34.8 16.7 37.8 25.0 16.4 55.6 16.0 9.5 15.4 21.4 14.3 31.5 35.9 40.8 22.7 40.9
Llama 3.1 8 33.3 34.5 34.1 21.7 16.7 40.5 45.0 27.3 22.2 30.0 28.6 38.5 14.3 28.6 38.9 33.3 48.7 45.5 42.4
Llama 3 8 28.6 34.5 34.1 26.1 16.7 40.5 40.0 23.6 33.3 26.0 23.8 38.5 0.0 14.3 40.7 41.0 40.1 31.8 43.9
DeepSeek-R1 8 23.8 42.9 36.4 30.4 16.7 35.1 20.0 27.3 22.2 30.0 23.8 61.5 21.4 0.0 33.3 25.6 34.2 45.5 37.9
GLM 4 9 47.6 56.0 50.0 39.1 33.3 21.6 20.0 25.5 11.1 22.0 19.0 30.8 14.3 14.3 27.8 23.1 43.4 54.5 56.1
Gemma 2 9 42.9 52.4 20.5 56.5 33.3 27.0 20.0 27.3 11.1 28.0 9.5 46.2 21.4 57.1 20.4 17.9 36.8 40.9 42.4
Mistral-Nemo 12 28.6 44.0 34.1 17.4 16.7 24.3 30.0 30.9 22.2 30.0 28.6 23.1 35.7 42.9 29.6 15.4 40.8 27.3 36.4
StableLM 2 12 23.8 31.0 20.5 30.4 16.7 10.8 5.0 7.3 11.1 8.0 9.5 0.0 7.1 14.3 24.1 10.3 32.2 27.3 24.2
Llama 2 13 23.8 26.2 6.8 13.0 33.3 16.2 5.0 14.5 11.1 14.0 0.0 30.8 7.1 0.0 16.7 7.7 13.2 13.6 15.2
Phi 3 14 19.0 38.1 34.1 30.4 16.7 40.5 50.0 32.7 33.3 30.0 28.6 30.8 14.3 42.9 38.9 38.5 40.8 29.5 39.4
@@
Mistral Small 22 0.0 13.1 0.0 4.3 0.0 0.0 0.0 1.8 0.0 2.0 4.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.5
Gemma 2 27 4.8 15.5 6.8 13.0 0.0 13.5 15.0 5.5 11.1 6.0 4.8 7.7 7.1 0.0 9.3 12.8 3.3 6.8 0.0
Aya-23 35 0.0 11.9 9.1 0.0 0.0 2.7 0.0 3.6 0.0 4.0 0.0 15.4 0.0 0.0 1.9 0.0 0.0 4.5 4.5
Command-R 35 0.0 6.0 4.5 4.3 0.0 2.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 2.6 0.7 4.5 0.0
Alfred 40 0.0 1.2 2.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 5.6 0.0 0.7 2.3 0.0
Mixtral 8x7 0.0 1.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 0.0 0.0 0.0 0.0
Llama 2 70 0.0 0.0 0.0 0.0 0.0 2.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 0.0 0.0 2.3 0.0
Llama 3 70 0.0 26.2 15.9 17.4 0.0 21.6 10.0 9.1 0.0 10.0 0.0 15.4 7.1 14.3 13.0 10.3 0.7 9.1 1.5
Llama 3.1 70 4.8 31.0 22.7 17.4 0.0 13.5 10.0 12.7 11.1 14.0 4.8 23.1 7.1 28.6 16.7 12.8 2.0 6.8 6.1
DeepSeek-R1 70 4.8 26.2 18.2 8.7 0.0 21.6 5.0 10.9 0.0 12.0 9.5 30.8 0.0 0.0 31.5 15.4 0.7 4.5 0.0
Qwen 2 72 0.0 13.1 11.4 8.7 0.0 5.4 15.0 7.3 0.0 8.0 4.8 7.7 14.3 28.6 5.6 10.3 2.6 11.4 3.0
Command-R+ 104 4.8 16.7 9.1 17.4 16.7 0.0 0.0 7.3 0.0 4.0 0.0 7.7 7.1 0.0 1.9 0.0 3.3 4.5 1.5
Mistral-Large 123 4.8 23.8 18.2 13.0 0.0 2.7 5.0 7.3 0.0 8.0 4.8 15.4 0.0 14.3 9.3 2.6 0.7 4.5 1.5
Mixtral 8x22 0.0 3.6 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.7 2.6 0.0 0.0 1.5
WizardLM 2 8x22 0.0 13.1 2.3 0.0 0.0 5.4 5.0 3.6 0.0 4.0 0.0 15.4 0.0 0.0 5.6 5.1 0.0 4.5 1.5
@@
Llama 3.1 405 42.9 67.9 68.2 47.8 33.3 37.8 40.0 41.8 22.2 38.0 33.3 53.8 28.6 14.3 55.6 33.3 55.9 63.6 59.1
DeepSeek-R1 671 81.0 94.0 93.2 91.3 100.0 70.3 65.0 69.1 55.6 70.0 57.1 84.6 71.4 100.0 81.5 64.1 79.6 86.4 84.8
ChatGPT 3.5 unk 71.4 91.7 90.9 100.0 50.0 73.0 60.0 74.5 44.4 76.0 71.4 92.3 71.4 71.4 75.9 69.2 85.5 88.6 83.3
Claude 3.5 S unk 42.9 85.7 88.6 100.0 50.0 70.3 75.0 89.1 66.7 90.0 81.0 100.0 92.9 85.7 85.2 69.2 88.2 86.4 86.4
ChatGPT 4o unk 52.4 86.9 90.9 95.7 66.7 78.4 60.0 74.5 55.6 76.0 66.7 92.3 78.6 85.7 74.1 64.1 83.6 86.4 83.3
ChatGPT 4o m unk 38.1 79.8 68.2 87.0 16.7 56.8 65.0 52.7 22.2 54.0 47.6 76.9 42.9 71.4 64.8 51.3 61.8 86.4 81.8
Gemini 2.0 F unk 33.3 81.0 84.1 87.0 16.7 70.3 70.0 72.7 66.7 76.0 66.7 84.6 71.4 85.7 68.5 53.8 80.9 86.4 84.8
LeChat unk 38.1 65.5 61.4 47.8 50.0 24.3 50.0 34.5 44.4 34.0 28.6 38.5 35.7 42.9 35.2 35.9 39.5 40.9 47.0
@@ G Fine-grained Category Description
H Fine-grained Category Annotation Prompt
We use the following prompt to categorise the exam questions:
I have a language test for Luxembourgish where I would like you to categorise every question
into different types of skill they are testing. For every question and every category,
determine and mark whether or not a given question belongs to a given category, as follows:
0, the question does not belong to the category
1, the question belongs to the category
Write the output in tabular form like TSV.
The file I am uploading are some questions that I already manually annotated.
Use it as a guideline.
The list of categories and explanations is as follows:
vocabulary: any question relating to vocabulary
grammar: any question relating to grammar
vocabulary_idiom: a question relating to idioms or metaphors
vocabulary_noun: a vocabulary question relating to a noun
vocabulary_verb: a vocabulary question relating to a verb
vocabulary_adjective: a vocabulary question relating to an adjective
vocabulary_native_word: a vocabulary question where the correct answer is a
word exclusive to Luxembourgish
grammar_gender: a question relating to determining the gender of a
determinant, adjective, or pronoun
grammar_pronouns: a question relating to pronouns
grammar_verbs:  a question relating to conjugations of verbs
grammar_adjectives:  a question relating to adjectives
grammar_tense: a question relating to the correct use of tenses
grammar_tense_past: a question relating to the past tense
grammar_tense_present: a question relating to the present tense
grammar_verbs_modal_verbs: a question relating to modal verbs
of prepositions or conjunctions
reading_comprehension_true_false: a reading comprehension question
where the answer is either true or false
reading_comprehension_simple: a reading comprehension question that
does not require to understand the entire text excerpt to answer correctly
reading_comprehension_complex: a reading comprehension question that
requires a deeper understanding of the entire text excerpt to answer correctly
I LuxGen Evaluation Results

### v4

@@ I LuxGen Evaluation Results using LLM-as-a-Judge metrics
Qwen 2 0.5 14.3 25.0 22.7 39.1 16.7 13.5 15.0 22.2 0.0 24.5 20.0 30.8 28.6 14.3 22.2 20.5 42.8 0.0 1.5
Llama 3.2 1 14.3 26.2 20.5 17.4 33.3 21.6 10.0 22.2 11.1 22.4 25.0 30.8 21.4 0.0 31.5 17.9 26.3 13.6 18.2
Gemma 3 1 33.3 32.1 18.2 21.7 33.3 35.1 20.0 38.9 55.6 34.7 40.0 46.2 21.4 28.6 35.2 35.9 58.6 29.5 34.8
Qwen 2 1.5 33.3 31.0 27.3 30.4 16.7 32.4 50.0 20.4 22.2 20.4 30.0 23.1 0.0 0.0 33.3 30.8 38.2 9.1 13.6
StableLM 2 1.6 19.0 16.7 22.7 17.4 16.7 13.5 10.0 24.1 11.1 26.5 20.0 23.1 28.6 28.6 27.8 12.8 10.5 4.5 9.1
EuroLLM 1.7 19.0 25.0 13.6 43.5 16.7 13.5 15.0 18.5 11.1 16.3 25.0 15.4 7.1 0.0 33.3 15.4 25.7 11.4 9.1
Gemma 2 2 33.3 35.7 27.3 43.5 33.3 37.8 35.0 31.5 22.2 32.7 40.0 23.1 35.7 0.0 25.9 23.1 44.7 18.2 21.2
Llama 3.2 3 28.6 34.5 27.3 34.8 33.3 18.9 15.0 27.8 44.4 24.5 20.0 38.5 21.4 14.3 35.2 20.5 39.5 22.7 30.3
Phi 3 3.8 23.8 25.0 36.4 21.7 33.3 27.0 35.0 24.1 11.1 26.5 25.0 30.8 14.3 42.9 25.9 20.5 40.1 29.5 33.3
Phi 3.5 3.8 9.5 17.9 11.4 26.1 0.0 16.2 20.0 24.1 0.0 22.4 20.0 30.8 14.3 42.9 18.5 12.8 20.4 20.5 16.7
Gemma 3 4 33.3 51.2 50.0 30.4 50.0 35.1 20.0 29.6 22.2 28.6 25.0 46.2 7.1 14.3 33.3 30.8 44.1 61.4 40.9
Qwen 2 7 33.3 38.1 34.1 30.4 50.0 37.8 40.0 25.9 33.3 28.6 25.0 30.8 21.4 42.9 25.9 41.0 45.4 40.9 37.9
Llama 2 7 28.6 32.1 27.3 34.8 33.3 37.8 20.0 22.2 33.3 16.3 15.0 38.5 0.0 0.0 27.8 28.2 46.7 22.7 33.3
WizardLM 2 7 19.0 33.3 40.9 34.8 16.7 40.5 30.0 25.9 22.2 26.5 25.0 15.4 50.0 42.9 29.6 41.0 35.5 15.9 34.8
Mistral 7 33.3 38.1 25.0 21.7 16.7 27.0 35.0 27.8 22.2 24.5 35.0 15.4 14.3 28.6 24.1 30.8 42.8 25.0 37.9
Aya-23 8 47.6 34.5 27.3 34.8 16.7 37.8 30.0 16.7 55.6 14.3 5.0 15.4 21.4 14.3 33.3 38.5 40.1 22.7 40.9
Llama 3.1 8 33.3 29.8 34.1 21.7 16.7 37.8 40.0 25.9 22.2 28.6 30.0 38.5 7.1 28.6 35.2 30.8 45.4 45.5 42.4
Llama 3 8 28.6 35.7 34.1 26.1 16.7 32.4 35.0 22.2 33.3 24.5 25.0 38.5 0.0 14.3 38.9 33.3 42.1 31.8 43.9
DeepSeek-R1 8 33.3 44.0 31.8 52.2 0.0 27.0 35.0 40.7 11.1 40.8 35.0 38.5 42.9 28.6 38.9 20.5 42.1 43.2 45.5
GLM 4 9 57.1 56.0 52.3 43.5 33.3 24.3 20.0 33.3 11.1 30.6 30.0 30.8 28.6 14.3 25.9 25.6 46.1 54.5 56.1
EuroLLM 9 38.1 42.9 27.3 60.9 50.0 32.4 50.0 29.6 22.2 26.5 25.0 30.8 7.1 28.6 33.3 33.3 40.8 38.6 40.9
Gemma 2 9 57.1 61.9 50.0 65.2 33.3 29.7 30.0 37.0 11.1 40.8 20.0 53.8 42.9 71.4 35.2 23.1 44.7 47.7 45.5
Mistral-Nemo 12 28.6 42.9 31.8 17.4 16.7 24.3 30.0 29.6 22.2 28.6 25.0 23.1 35.7 42.9 29.6 15.4 40.8 27.3 36.4
StableLM 2 12 23.8 28.6 25.0 30.4 16.7 8.1 5.0 3.7 11.1 4.1 5.0 0.0 0.0 14.3 25.9 7.7 30.9 27.3 24.2
Gemma 3 12 47.6 65.5 56.8 73.9 33.3 29.7 25.0 25.9 33.3 24.5 30.0 23.1 21.4 28.6 42.6 20.5 53.9 77.3 60.6
Llama 2 13 38.1 32.1 25.0 13.0 50.0 35.1 10.0 18.5 44.4 20.4 20.0 30.8 7.1 0.0 33.3 23.1 23.7 25.0 24.2
Phi 3 14 19.0 35.7 34.1 30.4 16.7 35.1 50.0 29.6 33.3 26.5 20.0 30.8 14.3 42.9 40.7 33.3 40.8 29.5 39.4
Phi 4 14 23.8 48.8 45.5 47.8 33.3 35.1 50.0 42.6 11.1 42.9 30.0 61.5 35.7 57.1 38.9 30.8 47.4 59.1 51.5
@@
Mistral Small 22 28.6 50.0 36.4 26.1 16.7 16.2 25.0 35.2 33.3 38.8 35.0 46.2 42.9 42.9 27.8 15.4 48.0 38.6 50.0
Gemma 2 27 47.6 66.7 52.3 60.9 50.0 35.1 45.0 29.6 11.1 28.6 25.0 30.8 21.4 42.9 46.3 35.9 51.3 77.3 72.7
Gemma 3 27 38.1 75.0 56.8 87.0 50.0 54.1 55.0 50.0 44.4 51.0 55.0 38.5 35.7 57.1 61.1 46.2 61.8 70.5 69.7
QWQ 32 47.6 66.7 45.5 60.9 33.3 40.5 40.0 42.6 11.1 42.9 45.0 38.5 42.9 71.4 35.2 35.9 40.8 70.5 63.6
Aya-23 35 28.6 44.0 31.8 47.8 16.7 32.4 30.0 29.6 22.2 26.5 25.0 23.1 21.4 42.9 31.5 30.8 39.5 43.2 50.0
Command-R 35 38.1 44.0 31.8 39.1 16.7 32.4 35.0 33.3 44.4 36.7 25.0 69.2 21.4 28.6 35.2 35.9 40.1 34.1 53.0
Alfred 40 23.8 32.1 38.6 34.8 33.3 29.7 20.0 37.0 11.1 36.7 35.0 53.8 21.4 28.6 25.9 35.9 44.7 36.4 30.3
Mixtral 8x7 28.6 46.4 50.0 43.5 16.7 27.0 30.0 24.1 22.2 26.5 25.0 38.5 14.3 42.9 35.2 23.1 37.5 34.1 51.5
Llama 2 70 33.3 32.1 31.8 17.4 16.7 27.0 20.0 24.1 33.3 24.5 20.0 46.2 7.1 0.0 33.3 25.6 48.0 34.1 39.4
Llama 3 70 52.4 58.3 45.5 56.5 33.3 35.1 25.0 31.5 33.3 30.6 25.0 38.5 14.3 28.6 37.0 33.3 42.8 50.0 48.5
Llama 3.1 70 28.6 63.1 56.8 65.2 33.3 35.1 35.0 48.1 44.4 46.9 40.0 38.5 57.1 57.1 53.7 30.8 47.4 52.3 54.5
DeepSeek-R1 70 47.6 73.8 56.8 47.8 16.7 29.7 40.0 46.3 11.1 49.0 55.0 38.5 42.9 42.9 51.9 17.9 46.1 65.9 66.7
Qwen 2 72 33.3 56.0 50.0 47.8 0.0 40.5 35.0 38.9 33.3 38.8 30.0 38.5 57.1 57.1 35.2 33.3 45.4 54.5 60.6
Command-R+ 104 23.8 47.6 43.2 52.2 50.0 21.6 0.0 38.9 11.1 36.7 30.0 46.2 35.7 42.9 44.4 5.1 46.1 43.2 40.9
Mistral-Large 123 42.9 75.0 56.8 56.5 50.0 45.9 50.0 40.7 33.3 40.8 35.0 61.5 21.4 14.3 37.0 35.9 44.1 68.2 50.0
Mixtral 8x22 38.1 48.8 54.5 56.5 33.3 54.1 45.0 38.9 44.4 38.8 35.0 53.8 21.4 42.9 35.2 43.6 45.4 52.3 54.5
WizardLM 2 8x22 28.6 47.6 36.4 26.1 16.7 29.7 35.0 37.0 33.3 34.7 40.0 53.8 14.3 0.0 48.1 23.1 46.1 52.3 51.5
@@
Llama 3.1 405 42.9 67.9 68.2 47.8 33.3 37.8 40.0 40.7 22.2 36.7 30.0 53.8 28.6 14.3 55.6 33.3 55.9 63.6 59.1
DeepSeek-R1 671 81.0 94.0 93.2 91.3 100.0 70.3 65.0 68.5 55.6 69.4 55.0 84.6 71.4 100.0 81.5 64.1 79.6 86.4 84.8
ChatGPT 3.5 unk 71.4 91.7 90.9 100.0 50.0 73.0 60.0 75.9 44.4 77.6 75.0 92.3 71.4 71.4 75.9 69.2 85.5 88.6 83.3
Claude 3.5 S unk 42.9 85.7 88.6 100.0 50.0 70.3 75.0 90.7 66.7 91.8 85.0 100.0 92.9 85.7 85.2 69.2 88.2 86.4 86.4
ChatGPT 4o unk 52.4 86.9 90.9 95.7 66.7 78.4 60.0 75.9 55.6 77.6 70.0 92.3 78.6 85.7 74.1 64.1 83.6 86.4 83.3
ChatGPT 4o m unk 38.1 79.8 68.2 87.0 16.7 56.8 65.0 53.7 22.2 55.1 50.0 76.9 42.9 71.4 64.8 51.3 61.8 86.4 81.8
Gemini 2.0 F unk 33.3 81.0 84.1 87.0 16.7 70.3 70.0 75.9 66.7 79.6 75.0 84.6 71.4 85.7 68.5 53.8 80.9 86.4 84.8
LeChat unk 38.1 65.5 61.4 47.8 50.0 24.3 50.0 35.2 44.4 34.7 30.0 38.5 35.7 42.9 35.2 35.9 39.5 40.9 47.0
@@
J Comparison of LLM-as-a-Judge and traditional metrics
In this section, we compare our LLM-as-a-Judge metrics to the BERTScore and METEOR scores in order to assess the usefulness of our novel metrics. Figure 3 and 4 show performances of every studied LLM on the Headline Generation and Short Description tasks using the four different metrics, respectively. We observe that while the scores are different for each metric they appear to be highly correlated to each other. This is further corroborated by Tables 12(a) through 13(c) which contain correlation matrices with Pearson Correlation Coefficients, t-scores, and p-values between every metric, showing strong correlations between the traditional metrics and the LLM-as-a-Judge metrics, with the significance tests showing the correlations to be significant at a significance level $\alpha=0.01$.
Figure 3: Comparison between LLM-as-a-Judge metrics, BERTScore, and METEOR on the headline generation task. We are using a stacked line chart to better highlight the correlations between the metrics.
Figure 4: Comparison between LLM-as-a-Judge metrics, BERTScore, and METEOR on the short description task. We are using a stacked line chart to better highlight the correlations between the metrics.
Claude-G Claude-A METEOR BERTScore
Claude-G 1.00 0.86 0.71 0.70
Claude-A 0.86 1.00 0.56 0.55
METEOR 0.71 0.56 1.00 0.95
BERTScore 0.70 0.55 0.95 1.00
(a) Pearson Correlation Coefficient
Claude-G Claude-A METEOR BERTScore
Claude-G / 11.90 7.27 7.07
Claude-A 11.90 / 4.82 4.72
METEOR 7.27 4.82 / 22.60
BERTScore 7.07 4.72 22.60 /
(b) T-Scores
Claude-G Claude-A METEOR BERTScore
Claude-G / 2.472E-16 2.050E-09 4.241E-09
Claude-A 2.472E-16 / 1.322E-05 1.859E-05
METEOR 2.051E-09 1.322E-05 / 3.134E-28
BERTScore 4.241E-09 1.859E-05 3.134E-28 /
(c) P-Values
Table 12: Correlation tests for different metrics on the headline generation task.
Claude-G Claude-A METEOR BERTScore
Claude-G 1.00 0.90 0.65 0.72
Claude-A 0.90 1.00 0.60 0.73
METEOR 0.65 0.60 1.00 0.89
BERTScore 0.72 0.73 0.89 1.00
(a) Pearson Correlation Coefficient
Claude-G Claude-A METEOR BERTScore
Claude-G / 15.08 6.13 7.36
Claude-A 15.08 / 5.33 7.67
METEOR 6.13 5.33 / 14.31
BERTScore 7.36 7.67 14.31 /
(b) T-Scores
Claude-G Claude-A METEOR BERTScore
Claude-G / 1.941E-20 1.266E-07 1.466E-09
Claude-A 1.941E-20 / 2.215E-06 4.740E-10
METEOR 1.266E-07 2.215E-06 / 1.736E-19
BERTScore 1.466E-09 4.740E-10 1.736E-19 /
(c) P-Values
Table 13: Correlation tests for different metrics on the Short Description task.
K Results for the Short Description Task
Figure 5 shows the LLM performances on the Short Description task plotted against their performances on the proficiency exams while Table 14 shows PCCs, t-scores, and corresponding p-values.
Task Small Medium Large All
Claude-G 0.670 0.773 0.926 0.779
Claude-A 0.700 0.812 0.859 0.660
METEOR 0.743 0.550 0.780 0.702
BERTScore 0.749 0.595 0.618 0.604
(a) Pearson Correlation Coefficients
small medium large all
Claude-G 6.443 8.702 17.516 8.884
Claude-A 7.005 9.935 11.970 6.279
METEOR 7.935 4.700 8.911 7.036
BERTScore 8.072 5.293 5.612 5.419
(b) T-Scores
small medium large all
Claude-G 4.126E-08 1.182E-11 3.185E-23 6.204E-12
Claude-A 5.338E-09 1.609E-13 1.980E-16 7.483E-08
METEOR 1.837E-10 2.009E-05 5.635E-12 4.770E-09
BERTScore 1.119E-10 2.563E-06 8.25E-07 1.64E-06
(c) P-Values
Table 14: Correlation tests between the overall exam scores and Short Description task performances.
(a) Claude Score - Grammar and Spelling
(b) Claude Score - Adequacy
(c) BERTScore
(d) METEOR Score
Figure 5: Performance on Language Exams vs Performance on the Short Description Task.